# Instance-Adaptive Keypoint Learning with Local-to-Global Geometric Aggregation for Category-Level Object Pose Estimation

### v1

@@ Abstract
Category-level object pose estimation aims to predict the 6D pose and size of previously unseen instances from predefined categories, requiring strong generalization across diverse object instances. Although many previous methods attempt to mitigate intra-class variations, they often struggle with instances exhibiting complex geometries or significant deviations from canonical shapes. To address this challenge, we propose INKL-Pose, a novel category-level object pose estimation framework that enables INstance-adaptive Keypoint Learning with local-to-global geometric aggregation. Specifically, our approach first predicts semantically consistent and geometric informative keypoints through an Instance-Adaptive Keypoint Generator, then refines them with: (1) a Local Keypoint Feature Aggregator capturing fine-grained geometries, and (2) a Global Keypoint Feature Aggregator using bidirectional Mamba for structural consistency. To enable bidirectional modeling in Mamba, we introduce a Feature Sequence Flipping strategy that preserves spatial coherence while constructing backward feature sequences. Additionally, we design a surface loss and a separation loss to enforce uniform coverage and spatial diversity in keypoint distribution. The generated keypoints are finally mapped to a canonical space for regressing the object’s 6D pose and size. Extensive experiments on CAMERA25, REAL275, and HouseCat6D demonstrate that INKL-Pose achieves state-of-the-art performance and significantly outperforms existing methods.
@@ I Introduction
To improve the quality of NOCS reconstructions, a common approach is to incorporate category-level shape priors. For example, SPD [11] learns object-specific deformations from predefined category templates and establishes dense correspondences between the observed point cloud and its reconstructed canonical shape. This paradigm has inspired a series of follow-up works [15, 16, 18, 19], which typically employ average shape point clouds as priors and design dedicated deformation networks to better adapt to intra-class variations. However, such priors remain fixed during both training and inference, limiting the model’s adaptability to diverse and unseen geometries. To address this, Query6DoF [20] introduces class-specific learnable query embeddings that are jointly optimized with the pose estimator, allowing the priors to adapt dynamically during training. More recently, shape-prior-free methods [12, 21, 13, 22, 23, 14] have drawn increasing attention due to their enhanced generalization to novel object instances. Among them, AG-Pose [14] introduces a keypoint-based method that predicts representative keypoints for each instance and establishes keypoint-level correspondences, achieving notable improvements against structural discrepancies. Nevertheless, its keypoint features are derived solely from local attention within fixed neighborhoods and simple global average pooling, limiting their capacity to capture fine-grained geometric details and hindering the model’s ability to perceive global structural consistency. Moreover, the keypoint distribution is regulated by heuristic losses, lacking explicit constraints to ensure uniform surface coverage and sufficient spatial separation. As a result, the predicted keypoints (as illustrated in Fig. 1) often cluster in limited regions, fail to accurately capture the object’s surface coverage, and exhibit substantial NOCS errors, particularly on objects with complex geometries.
@@ II-B Category-Level Object Pose Estimation
To overcome this limitation, Query6DoF [20] introduces class-specific learnable query embeddings that are jointly optimized with the pose estimation network, enabling the shape priors to adapt dynamically during training. In addition, shape-prior-free methods [12, 21, 13, 22, 23, 14] have attracted increasing attention for their improved generalization to diverse and unseen object geometries. For instance, IST-Net [13] adopts an implicit spatial transformation framework that directly learns correspondences between camera-space and world-space features without relying on explicit priors. CLIPose [12] leverages a vision-language model to extract rich semantic features from both text and images, aligning them with point cloud representations via contrastive learning. TG-Pose [21] introduces topological descriptors and a symmetry-aware reconstruction module to refine object geometry and enhance pose stability. SecondPose [23] integrates geometric features with semantic embeddings from DINOv2 [28], supervised under SE(3)-consistent constraints to improve cross-instance generalization. SpherePose [29] employs a shared spherical proxy representation to address semantic inconsistency arising from shape-dependent canonical spaces. More recently, AG-Pose [14] proposes detecting instance-adaptive keypoints and establishing keypoint-level correspondences to enhance robustness against structural variations across object instances, achieving notable performance improvements.
@@ III-A Overview
Figure 2: (a) Overall framework of INKL-Pose. (b) The Instance-Adaptive Keypoint Generator dynamically generates keypoints along with corresponding features by assigning weights based on the semantic and geometric information extracted by the Pointwise Feature Encoder. (c) The Local Keypoint Feature Aggregator captures fine-grained local structures by aggregating point cloud neighborhood information. (d) The Global Keypoint Feature Aggregator models long-range dependencies using bidirectional Mamba, extracting globally consistent geometric representations across instances.
@@ III-C Instance-Adaptive Keypoint Generator
Specifically, we formulate keypoint generation as a learnable query-based process that adapts to both semantic and geometric variations, following the paradigm introduced in AG-Pose [14]. Unlike AG-Pose, however, our approach introduces two novel loss functions that explicitly encourage the predicted keypoints to be uniformly distributed and spatially diverse across the object surface. As shown in Fig. 2(b), we first initialize a learnable keypoint query embedding $\mathcal{Q}_{kpt}$, which is initially set to zero. Then, through multi-head cross-attention and multi-head self-attention mechanisms, the keypoint query vectors interact comprehensively with both the input features $\mathcal{F}_{o}$ and themselves, progressively updating to form instance-specific query embeddings. Finally, we compute the cosine similarity between $\mathcal{Q}_{kpt}$ and $\mathcal{F}_{o}$, and apply a softmax function to normalize the results, yielding the keypoint probability distribution matrix $\mathcal{W}$:
$\mathcal{W}=Softmax\left(\frac{\mathcal{Q}_{kpt}\cdot\mathcal{F}_{o}}{\|\mathcal{Q}_{kpt}\|_{2}\times\|\mathcal{F}_{o}\|_{2}+\varphi}\right),$ (1)
@@
$\mathcal{P}_{kpt}=\mathcal{W}\times\mathcal{P}_{o},\quad\mathcal{F}_{kpt}=\mathcal{W}\times\mathcal{F}_{o}.$ (2)
@@
Second, the separation loss $\mathcal{L}_{sep}$ promotes geometric diversity by encouraging each keypoint to be spatially distinct from its local neighbors. Specifically, it maximizes the average pairwise distance between each predicted keypoint and its $M$-nearest neighbors:
$\mathcal{L}_{sep}=\frac{1}{\max\left(\frac{1}{M}\sum_{i=1}^{M}\|\mathcal{P}_{kpt}^{i}-\text{NN}(\mathcal{P}_{kpt}^{i})\|_{2},\,\epsilon\right)},$ (4)
where $\text{NN}(\cdot)$ denotes the nearest neighbors in Euclidean space, and the denominator is lower-bounded by a small constant $\epsilon=0.01$ to ensure numerical stability. In our experiments, we set $M=2$, which provides a lightweight yet effective constraint that mitigates keypoint clustering while preserving flexibility in spatial distribution.
@@ III-D Local Keypoint Feature Aggregator
Next, for each keypoint, we construct a local neighborhood in both coordinate and feature spaces using the $K$-nearest neighbor ($K\!N\!N$) algorithm. This yields coordinate neighborhoods $\mathcal{P}_{K}\in\mathbb{R}^{N_{kpt}\times K\times 3}$ and feature neighborhoods $\mathcal{F}_{K}\in\mathbb{R}^{N_{kpt}\times K\times d}$, where $K$ is the neighborhood size. We then concatenate these two and apply standard normalization to obtain fused local group features $\mathcal{F}_{K}^{\prime}\in\mathbb{R}^{N_{kpt}\times K\times(d+3)}$, which jointly encode geometric structure and semantic information:
@@ III-E1 Preliminaries
SSMs [30] were originally introduced in control theory for modeling dynamical systems. The core principle is to map the input signal $x(t)$ to an output signal $y(t)$ through a latent state $h(t)$, whose evolution is governed by both the current state and the input. This process is described by the following system of ordinary differential equations:
@@
where $A$ is the state transition matrix, and $B$ and $C$ are the input and output projection matrices, respectively, all of which can be parameterized and learned within a neural network.
As this formulation is defined over continuous time, it must be discretized before it can be applied to discrete data such as point clouds. Mamba [31] achieves this using the zero-order hold method, which introduces a time step $\Delta$ to convert the continuous parameters $A$ and $B$ into discrete versions $\overline{A}$ and $\overline{B}$. Here, $\Delta$ reflects the temporal resolution of the sampled input sequence:
$\overline{A}=e^{\Delta A},\quad\overline{A}=(\Delta A)^{-1}(e^{\Delta A}-I)\Delta B.$ (10)
With the discretized parameters, Eq. (9) can be reformulated into a recurrent sequence model:
@@
where $t$ denotes the time step. This recurrent formulation enables sequence-to-sequence mapping from inputs $x_{t}$ to outputs $y_{t}$, but its inherently sequential nature poses challenges for efficient training.
@@ III-E2 GKFA Module
To capture global semantic consistency and long-range dependencies among keypoints, we integrate a bidirectional variant of Mamba [34] into the GKFA module. This design enables bidirectional information flow across keypoints, facilitating the learning of globally consistent geometric structures. Traditionally, bidirectional modeling is achieved through a strategy we refer to as Point Sequence Flipping (PSF), where the input point sequence is reversed based on spatial indices. However, such input-order-based reversal is not well suited for point cloud data, as the unordered and non-causal nature of point sets may disrupt spatial and semantic coherence, especially when geometrically unrelated points are juxtaposed in the reversed sequence.
@@
$\mathcal{L}_{sim}(\mathcal{P}_{o},\mathcal{P}_{rec})=$ $\,\frac{1}{N_{pts}}\sum_{x_{i}\in\mathcal{P}_{o}}\min_{y_{j}\in\mathcal{P}_{rec}}\|x_{i}-y_{j}\|_{2}^{2}$ (16)
$+\frac{1}{N_{rec}}\sum_{y_{j}\in\mathcal{P}_{rec}}\min_{x_{i}\in\mathcal{P}_{o}}\|y_{j}-x_{i}\|_{2}^{2}.$
@@
To further strengthen the representational capacity of keypoints, we stack multiple LKFA and GKFA modules in a cascaded architecture. Each block refines keypoint features by alternately aggregating fine-grained local geometry and long-range global context. This design allows hierarchical learning across different scales, improving the network’s ability to model complex object structures. As demonstrated by our ablation studies (Section IV-C), using 12 stacked blocks yields the best performance.
@@ III-F Pose and Size Estimator
$\mathcal{F}_{pose}=Concat[\mathcal{P}_{kpt},\mathcal{P}_{kpt}^{nocs},\mathcal{F}_{kpt}],$ (19)
@@
where $R_{pre}$ and $t_{pre}$ represent the predicted 6D object pose, and $s_{pre}$ represents the predicted object size. We simply use the $\mathcal{L}_{1}$ loss function to supervise $R_{pre}$, $t_{pre}$ and $s_{pre}$:
@@ III-G Overall Loss Function
$\mathcal{L}_{all}$ $=\omega_{sep}\mathcal{L}_{sep}+\omega_{surf}\mathcal{L}_{surf}$ (22)
$+\omega_{sim}\mathcal{L}_{sim}+\omega_{map}\mathcal{L}_{map}+\omega_{pose}\mathcal{L}_{pose},$
where $\omega_{sep}$, $\omega_{far}$, $\omega_{sim}$, $\omega_{map}$, and $\omega_{pose}$ are hyperparameters used to balance the contributions of separation loss, surface loss, similarity loss, mapping loss and pose loss.
@@ IV-C Ablation Studies
Figure 7: Performance comparison between INKL-Pose and state-of-the-art methods on the REAL275 and HouseCat6D real-world datasets.
@@ V Conclusion
In this paper, we proposed INKL-Pose, a novel keypoint-based framework for category-level 6D object pose estimation that enhances generalization to unseen object instances through instance-adaptive keypoint learning and local-to-global geometric aggregation. Specifically, we introduce an Instance-Adaptive Keypoint Generator that predicts semantically consistent and geometrically informative keypoints, along with two complementary modules for feature refinement: a Local Aggregator that captures fine-grained geometric cues and a Global Aggregator based on bidirectional Mamba with a Feature Sequence Flipping strategy for structure-aware global perception. Furthermore, we design surface loss and separation loss to explicitly supervise keypoint distribution, providing strong geometric guidance for robust pose estimation. Extensive experiments on CAMERA25, REAL275, and HouseCat6D demonstrate that INKL-Pose achieves state-of-the-art performance across both synthetic and real-world datasets, significantly outperforming existing approaches. In future work, we will focus on extending our framework to dynamic and temporal settings, enabling sequential pose tracking, and explore uncertainty-aware keypoint modeling to further improve robustness under occlusion and sensor noise.

### v2

@@ Abstract
Category-level object pose estimation aims to predict the 6D pose and size of previously unseen instances from predefined categories, requiring strong generalization across diverse object instances. Although many previous methods attempt to mitigate intra-class variations, they often struggle with instances exhibiting complex geometries or significant deviations from canonical shapes. To address this challenge, we propose INKL-Pose, a novel category-level object pose estimation framework that enables INstance-adaptive Keypoint Learning with local-to-global geometric aggregation. Specifically, our approach first predicts semantically consistent and geometrically informative keypoints through an Instance-Adaptive Keypoint Generator, then refines them with: (1) a Local Keypoint Feature Aggregator capturing fine-grained geometries, and (2) a Global Keypoint Feature Aggregator using bidirectional Mamba for structural consistency. To enable bidirectional modeling in Mamba, we introduce a Feature Sequence Flipping strategy that preserves spatial coherence while constructing backward feature sequence. Additionally, we design a surface loss and a separation loss to enforce uniform coverage and spatial diversity in keypoint distribution. The generated keypoints are finally mapped to a canonical space for regressing the object’s 6D pose and size. Extensive experiments on CAMERA25, REAL275, and HouseCat6D demonstrate that INKL-Pose achieves state-of-the-art performance and significantly outperforms existing methods.
@@ I Introduction
To improve the quality of NOCS reconstructions, a common approach is to incorporate category-level shape priors. For example, SPD [11] learns object-specific deformations from predefined category templates and establishes dense correspondences between the observed point cloud and its reconstructed canonical shape. This paradigm has inspired a series of follow-up works [15, 16, 18, 19], which typically employ average shape point clouds as priors and design dedicated deformation networks to better adapt to intra-class variations. However, such priors remain static during both training and inference, limiting the model’s adaptability to diverse and unseen geometries. To address this, Query6DoF [20] introduces class-specific learnable query embeddings that are jointly optimized with the pose estimator, allowing the priors to adapt dynamically during training. More recently, shape-prior-free methods [12, 21, 13, 22, 23, 14] have drawn increasing attention due to their enhanced generalization to novel object instances. Among them, AG-Pose [14] introduces a keypoint-based method that predicts representative keypoints for each instance and establishes keypoint-level correspondences, achieving notable improvements against structural discrepancies. Nevertheless, its keypoint features are derived solely from local attention within fixed neighborhoods and simple global average pooling, limiting their capacity to capture fine-grained geometric details and hindering the model’s ability to perceive global structural consistency. Moreover, the keypoint distribution is regulated by heuristic losses, lacking explicit constraints to ensure uniform surface coverage and sufficient spatial separation. As a result, the predicted keypoints (as illustrated in Fig. 1) often cluster in limited regions, fail to accurately capture the object’s surface coverage, and exhibit substantial NOCS errors, particularly on objects with complex geometries.
@@ II-B Category-Level Object Pose Estimation
To overcome this limitation, Query6DoF [20] introduces class-specific learnable queries that are jointly optimized with the pose estimation network, enabling the shape priors to adapt dynamically during training. In addition, shape-prior-free methods [12, 21, 13, 22, 23, 14] have attracted increasing attention for their improved generalization to diverse and unseen object geometries. IST-Net [13] adopts an implicit spatial transformation framework that directly learns correspondences between camera-space and world-space features without relying on explicit priors. CLIPose [12] leverages a vision-language model to extract rich semantic features from both text and images, aligning them with point cloud representations via contrastive learning. TG-Pose [21] introduces topological descriptors and a symmetry-aware reconstruction module to refine object geometry and enhance pose stability. SecondPose [23] integrates geometric features with semantic embeddings from DINOv2 [28], supervised under SE(3)-consistent constraints to improve cross-instance generalization. SpherePose [29] employs a shared spherical proxy representation to address semantic inconsistency arising from shape-dependent canonical spaces. CleanPose [30] integrates causal learning with knowledge distillation to address the spurious correlations introduced by confounding factors in pose estimation. More recently, AG-Pose [14] proposes detecting instance-adaptive keypoints and establishing keypoint-level correspondences to enhance robustness against structural variations across object instances, achieving notable performance improvements.
@@ III-A Overview
Figure 2: (a) Overall framework of INKL-Pose [^1] This figure was inspired by the drawing style of CleanPose [30]. We gratefully acknowledge the authors for their work.. (b) The Instance-Adaptive Keypoint Generator dynamically generates keypoints along with corresponding features by assigning weights based on the semantic and geometric information extracted by the Pointwise Feature Encoder. (c) The Local Keypoint Feature Aggregator captures fine-grained local structures by aggregating point cloud neighborhood information. (d) The Global Keypoint Feature Aggregator models long-range dependencies using bidirectional Mamba, extracting globally consistent geometric representations across instances.
@@ III-C Instance-Adaptive Keypoint Generator
Specifically, we formulate keypoint generation as a learnable query-based process that adapts to both semantic and geometric variations, following the paradigm introduced in AG-Pose [14]. In contrast to AG-Pose, our approach introduces two novel loss functions that explicitly encourage the predicted keypoints to be uniformly distributed and spatially diverse across the object surface. As shown in Fig. 1(b), we first initialize a learnable keypoint query embedding $\mathcal{Q}_{kpt}$, which is initially set to zero, providing a neutral starting point that avoids introducing bias toward specific patterns. This initialization facilitates stable learning dynamics and allows the query vectors to adapt purely based on the input data through the attention mechanism. Then, through multi-head cross-attention and multi-head self-attention mechanisms, the keypoint query vectors interact comprehensively with both the input features $\mathcal{F}_{o}$ and themselves, progressively updating to form instance-specific query embeddings. Finally, we compute the cosine similarity between $\mathcal{Q}_{kpt}$ and $\mathcal{F}_{o}$, and apply a softmax function to normalize the results, yielding the keypoint probability distribution matrix $\mathcal{W}$:
$\mathcal{W}=Softmax\left(\frac{\mathcal{Q}_{kpt}\cdot\mathcal{F}_{o}}{\|\mathcal{Q}_{kpt}\|_{2}\cdot\|\mathcal{F}_{o}\|_{2}+\varphi}\right),$ (1)
@@
$\mathcal{P}_{kpt}=\mathcal{W}\mathcal{P}_{o},\quad\mathcal{F}_{kpt}=\mathcal{W}\mathcal{F}_{o}.$ (2)
@@
Second, the separation loss $\mathcal{L}_{sep}$ promotes geometric diversity by encouraging each keypoint to be spatially distinct from its local neighbors. Specifically, it maximizes the average pairwise distance between each predicted keypoint and its $M$-nearest neighbors in Euclidean space:
$\mathcal{L}_{sep}=\frac{1}{\max\left(\frac{1}{N_{kpt}}\sum_{i=1}^{N_{kpt}}\frac{1}{M}\sum_{j=1}^{M}\|\mathcal{P}_{kpt}^{i}-\text{NN}_{j}(\mathcal{P}_{kpt}^{i})\|_{2},\epsilon\right)},$ (4)
where $\text{NN}_{j}(\mathcal{P}_{kpt}^{i})$ denotes the $j$th nearest neighbor of the $i$th keypoint, and $M$ is the number of neighbors considered. In our experiments, we set $M=2$, which provides a lightweight yet effective constraint to mitigate keypoint clustering while maintaining flexibility in spatial layout. The denominator is lower-bounded by a small constant $\epsilon$ = 0.01 to ensure numerical stability.
@@ III-D Local Keypoint Feature Aggregator
Next, for each keypoint, we construct a local neighborhood in both coordinate and feature spaces using the $K$-nearest neighbor ($K\!N\!N$) algorithm. This yields coordinate neighborhoods $\mathcal{P}_{K}\in\mathbb{R}^{N_{kpt}\times K\times 3}$ and feature neighborhoods $\mathcal{F}_{K}\in\mathbb{R}^{N_{kpt}\times K\times d}$, where $K$ is the neighborhood size. We then concatenate the spatial and feature descriptors and apply standard normalization to obtain fused local group features $\mathcal{F}_{K}^{\prime}\in\mathbb{R}^{N_{kpt}\times K\times(d+3)}$, which jointly encode geometric structure and semantic information:
@@ III-E1 Preliminaries
SSMs [31] were originally introduced in control theory for modeling linear time-invariant (LTI) dynamical systems. The core principle is to map the input signal $x(t)$ to an output signal $y(t)$ through a latent state $h(t)$, whose evolution is governed by both the current state and the input. This process is described by the following system of ordinary differential equations:
@@
where $A$ is the state transition matrix, and $B$ and $C$ are the input and output projection matrices, respectively. Here, $t$ denotes continuous time, and the system captures how the internal dynamics evolve smoothly over time.
To apply SSMs to discrete data such as point cloud sequences or tokens in deep learning, the continuous-time formulation must be discretized. Mamba [32] achieves this via the zero-order hold method, which introduces a time step ($\Delta$) to approximate the continuous dynamics. This leads to a transformation of the continuous matrices $A$ and $B$ into their discrete counterparts $\overline{A}$ and $\overline{B}$, as follows:
$\overline{A}=e^{\Delta A},\quad\overline{B}=(\Delta A)^{-1}(e^{\Delta A}-I)\Delta B.$ (10)
Here, $\Delta$ represents the temporal resolution of the sampled sequence (i.e., the interval between adjacent observations in discrete time). With these discretized parameters, the continuous-time system in Eq. (9) can be reformulated into a recurrent state-space model defined over discrete time steps:
@@
where $t$ denotes the discrete time step. This formulation enables efficient sequence-to-sequence mapping from inputs $x_{t}$ to outputs $y_{t}$, but its inherently sequential nature poses challenges for efficient training.
@@ III-E2 GKFA Module
As illustrated in Fig. 1(d), to capture global semantic consistency and long-range dependencies among keypoints, we integrate a bidirectional variant of Mamba [35] into the GKFA module. This design enables bidirectional information flow across keypoints, facilitating the learning of globally consistent geometric structures. Traditionally, bidirectional modeling is achieved through a strategy we refer to as Point Sequence Flipping (PSF), where the input point sequence is reversed based on spatial indices. However, such input-order-based reversal is not well suited for point cloud data, as the unordered and non-causal nature of point sets may disrupt spatial and semantic coherence, especially when geometrically unrelated points are juxtaposed in the reversed sequence.
@@
$\mathcal{L}_{sim}(\mathcal{P}_{o},\mathcal{P}_{rec})=$ $\,\frac{1}{N_{pts}}\sum_{x_{i}\in\mathcal{P}_{o}}\min_{y_{j}\in\mathcal{P}_{rec}}\|x_{i}-y_{j}\|_{2}^{2}+$ (16)
$\frac{1}{N_{rec}}\sum_{y_{j}\in\mathcal{P}_{rec}}\min_{x_{i}\in\mathcal{P}_{o}}\|y_{j}-x_{i}\|_{2}^{2}.$
@@
To further strengthen the representational capacity of keypoints, we stack multiple LKFA and GKFA modules in a cascaded architecture. Each block refines keypoint features by alternately aggregating fine-grained local geometry and long-range global context. This design allows hierarchical learning across different scales, improving the network’s ability to model complex object structures. As demonstrated by our ablation studies (Section IV-C), using $S$ = 12 stacked blocks yields the best performance.
@@ III-F Pose and Size Estimator
$\mathcal{F}_{pose}=Concat(\mathcal{P}_{kpt},\mathcal{P}_{kpt}^{nocs},\mathcal{F}_{kpt}),$ (19)
@@
where $R_{pre}$ and $t_{pre}$ represent the predicted 6D object pose, and $s_{pre}$ represents the predicted object size. We simply use the $\mathcal{L}_{2}$ norm loss function to supervise $R_{pre}$, $t_{pre}$ and $s_{pre}$:
@@ III-G Overall Loss Function
$\mathcal{L}_{all}=$ $\omega_{sep}\mathcal{L}_{sep}+\omega_{surf}\mathcal{L}_{surf}+$ (22)
$\omega_{sim}\mathcal{L}_{sim}+\omega_{map}\mathcal{L}_{map}+\omega_{pose}\mathcal{L}_{pose},$
where $\omega_{sep}$, $\omega_{surf}$, $\omega_{sim}$, $\omega_{map}$, and $\omega_{pose}$ are hyperparameters used to balance the contributions of separation loss, surface loss, similarity loss, mapping loss and pose loss.
@@ IV-C Ablation Studies
Figure 7: Performance comparison between INKL-Pose and state-of-the-art methods on the REAL275 and HouseCat6D real-world datasets [^2] This figure was inspired by the drawing style of CleanPose [30]. We gratefully acknowledge the authors for their work..
@@ V Conclusion
In this paper, we propose INKL-Pose, a novel keypoint-based framework for category-level object pose estimation that enhances generalization to unseen object instances through instance-adaptive keypoint learning and local-to-global geometric aggregation. Specifically, we introduce an Instance-Adaptive Keypoint Generator that predicts semantically consistent and geometrically informative keypoints, along with two complementary modules for feature refinement: a Local Keypoint Feature Aggregator that captures fine-grained geometric details and a Global Keypoint Feature Aggregator based on bidirectional Mamba with a Feature Sequence Flipping strategy for structure-aware global perception. Furthermore, we design a surface loss and a separation loss to explicitly supervise keypoint distribution, providing strong geometric guidance for robust pose estimation. Extensive experiments on CAMERA25, REAL275, and HouseCat6D demonstrate that INKL-Pose achieves state-of-the-art performance across both synthetic and real-world datasets, significantly outperforming existing approaches. In future work, we will focus on extending our framework to dynamic and temporal settings, enabling sequential pose tracking, and explore uncertainty-aware keypoint modeling to further improve robustness under occlusion and sensor noise.